Under standard conditions, what reaction occurs, if any, when each of the following operations are performed? For the reactions that occur, write a balanced, net ionic equation and calculate ℰ°, ΔG°, and K at 25°C. (Use the lowest possible coefficients. Include states-of-matter at 25°C and 1 atm in your answer. If an answer does not exist, enter DNE.)

a. Cl2 gas is bubbled into a solution of NaI.

______________________________

ℰ°     ____ V
ΔG°     ____  kJ
K    

b. An acidic solution of FeSO4 is exposed to air.

______________________________

ℰ°     ____  V
ΔG°     ____  kJ
K    
Answer #1

Part a.

Cl2(g) + 2NaI(aq) ----> I2(s) + 2NaCl(aq)

Net ionic equation: Cl2(g) + 2I-(aq) ----> I2(s) + 2Cl-(aq)

Eocell = EoCl2/Cl- - EoI2/I- = 1.36 V - 0.54 V = 0.82 V

The positive value of Eocell indicates that the cell reaction is spontaneous.

\DeltaGo = -nFEocell = -2*96500 C * 0.82 V = -158260 J or -158.26 kJ

\DeltaGo = -RT ln(K)

i.e. -158260 J/mol = -8.314 J/mol.K * 298 K * ln(K)

i.e. ln(K) = 63.877

i.e. K = e63.877 = 5.51*1027

Part b.

Net ionic equation: 2Fe2+ + 2H+ ----> 2Fe3+(aq) + H2 (DNE)

Eocell = EoFe2+/Fe3+ - EoH+/H2 = -0.77 V - 0 V = -0.77 V

The negative value of Eocell indicates that the cell reaction is nonspontaneous.

\DeltaGo = -nFEocell = -2*96500 C * (-0.77 V) = 148610 J or 148.61 kJ

\DeltaGo = -RT ln(K)

i.e. 148610 J/mol = -8.314 J/mol.K * 298 K * ln(K)

i.e. ln(K) = -59.982

i.e. K = e-59.982 = 8.91*10-27